# The Pell sequence and cyclotomic matrices involving squares over finite fields

### v2

@@ Abstract.
where $Q_{p}$ is the $p$-th term of the companion Pell sequence $\{Q_{i}\}_{i=0}^{\infty}$ defined by $Q_{0}=Q_{1}=2,Q_{i+1}=2Q_{i}+Q_{i-1}$.
@@ 1.2. Background and Motivation
[…]
@@
Vsemirnov [16, 17] later confirmed this tough conjecture by using ingenious and sophisticated matrix decomposition.
@@
[…]
@@
[…]
@@
$\left\{s_{1}=1,s_{2},\cdots,s_{n}\right\}=\{x^{2}:x\in\mathbb{F}_{q}^{\times}\}.$
Then we define the matrix
(1.5) $B_{q}(m):=\left[\left(s_{i}+s_{j}\right)^{m}\right]_{2\leq i,j\leq n},$
which concerns all nonzero squares over $\mathbb{F}_{q}$.
@@
(1.6) $\det\left[h(x_{i}+y_{j})\right]_{1\leq i,j\leq m}=a_{m-1}^{l}\cdot\prod_{r=0}^{l-1}\binom{l-1}{r}\cdot\prod_{1\leq i<j\leq m}\left(x_{i}-x_{j}\right)\left(y_{j}-y_{i}\right).$
By this we immediately see that $B_{q}(m)$ is singular whenever $m\leq n-3$. Hence it is meaningful to consider the cases of $n-2\leq m\leq q-1$. In this paper, we focus on the cases of $m=n-2,n-1,n$.
@@ 1.3. The Pell sequence
$Q_{i}=\left((1+\sqrt{2})^{i}+(1-\sqrt{2})^{i}\right).$
@@
1.4. The first result: case $m\in\{n-2,n-1\}$
Now we state our first result.
@@ 1.4. The first result: case $m\in\{n-2,n-1\}$
where $h(-p)$ is the class number of the imaginary quadratic field $\mathbb{Q}(\sqrt{-p})$.
[…]
@@
Remark 1.1.
With the help of a computer, we verify that
$\left\{p:\ p\ \text{is a prime with}\ 7\leq p\leq 10^{5}\ \text{and}\ Q_{p}\equiv 2\ ({\rm{mod}}\ p^{2}\mathbb{Z})\right\}=\left\{13,31\right\}.$
@@
1.5. The second result: case $m=n$
@@
$b_{p}=\frac{2p+(\frac{2}{p})-P_{p}}{2p}\ {\rm mod}\ p\mathbb{Z}\in\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}.$
@@
$P_{p}\equiv 2p+\left(\frac{2}{p}\right)\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
1.6. The third result: a variant of Carlitz’s cyclotomic matrices
@@
$G_{q}(\phi)=\begin{cases}(-1)^{f-1}\sqrt{q}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ (-1)^{f-1}{\bf i}^{f}\sqrt{q}&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4).\\ \end{cases}$
@@
[…]
@@
Recall that $n=(q-1)/2$ and $s_{1}=1,s_{2},\cdots,s_{n}$ are exactly all the non-zero squares over $\mathbb{F}_{q}$. Also, for any integers $a,b$ with $a\leq b$, we use the symbol $[a,b]$ to denote the set $\{a,a+1,\cdots,b\}$.
@@
Recall that the Pell sequence $\{P_{i}\}_{i=0}^{\infty}$ is defined by
$P_{0}=0,P_{1}=1,P_{i+1}=2P_{i}+P_{i-1},$
and its companion $\{Q_{i}\}_{i=0}^{\infty}$ is defined by
$Q_{0}=Q_{1}=2,Q_{i+1}=2Q_{i}+Q_{i-1}.$
We now need the following results, which were obtained by Z.-H. Sun [10, Theorem 4.1], Z.-W. Sun [11, Final Remark] and Z.-W. Sun [12, Remark 3.1] respectively.
@@
where $B_{p-3}$ is the $(p-3)$-th Bernoulli number.
@@
$F(t)=t^{n-1}+t^{n-2}+\cdots+t+1=\prod_{2\leq j\leq n}\left(t-s_{j}\right)=\sum_{k=0}^{n-1}(-1)^{k}\sigma_{k}\left(s_{2},\cdots,s_{n}\right)t^{n-k},$
@@
Suppose first that $f\geq 2$. As $q=p^{f}\geq 9$, we have
@@
$\det B_{q}(n-2)$ $=\prod_{r=0}^{n-2}\binom{n-2}{r}\cdot\prod_{2\leq i<j\leq n}-\left(s_{j}-s_{i}\right)^{2}$
@@
[…]
@@
where $h(-p)$ is the class number of the imaginary quadratic field $\mathbb{Q}(\sqrt{-p})$. This completes the proof of (i).
[…]
$-1\equiv(p-1)!\equiv(-1)^{n}\cdot(n!)^{2}\equiv(n!)^{2}\ ({\rm{mod}}\ p\mathbb{Z}).$
@@
where the last equality follows from $n(n-1)=3/4$ over $\mathbb{F}_{p}$.
[…]
@@ 3.2. Some $p$-adic preparations
Recall that $q=p^{f}$. Let $\zeta_{q-1}\in\mathbb{C}_{p}$ be a primitive $(q-1)$-th root of unity. Then it is known that $\mathbb{Q}_{p}(\zeta_{q-1})/\mathbb{Q}_{p}$ is an unramified extension and $[\mathbb{Q}_{p}(\zeta_{q-1}):\mathbb{Q}_{p}]=f$. Hence
@@
As $10\leq r\leq n-$ 1, we have $x_{0}(r)<1/2$. Hence by Eq. (3.3) we obtain
@@
Recall that $p$ is an odd prime. We next introduce the $p$-adic Gamma function $\Gamma_{p}$. For any $n\in\mathbb{Z}^{+}$, we define
$\Gamma_{p}(n)=(-1)^{n}\prod_{0<k<n,p\nmid k}k\in\mathbb{Z}_{p}^{\times}.$
@@
Let notations be as in Section 3. In this section, we fix a generator $g$ of the cyclic group $\mathbb{F}_{q}^{\times}$. Recall that $n=(q-1)/2$. For any $i\in[0,n-1]$, we let
@@
(4.2) $\det B_{q}(n)=\det\left[\omega_{q}\left(1+g^{2j-2i}\right)\right]_{1\leq i,j\leq n-1}\ {\rm mod}\ \mathfrak{p}=\det W_{n}({\bm{v}})\ {\rm mod}\ \mathfrak{p},$
@@
Note first that in this case, $\mathbb{Q}_{p}(\zeta_{q-1})=\mathbb{Q}_{p}$ and $\mathfrak{p}=p\mathbb{Z}_{p}$. By Eq. (4.5), for any $r\in[1,n-1]$, we see that
@@
We next consider $\lambda_{r}\ {\rm mod}\ p\mathbb{Z}_{p}$. By Lemma 3.5 and Lemma 3.4. Suppose $r\in[1,n-1]$. Then
$\lambda_{r}$ $\equiv\frac{(-1)^{r}}{2}J_{q}\left(\omega_{p}^{-n},\omega_{p}^{-r}\right)$
@@
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\prod_{r=0}^{n}\binom{n}{r}\cdot\left(1-2\sum_{r=0}^{n}\binom{n}{r}^{-1}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(1-\frac{n+1}{2^{n}}\sum_{k=1}^{n+1}\frac{2^{k}}{k}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(1-\frac{1}{2^{n+1}}\sum_{k=1}^{n+1}\frac{2^{k}}{k}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(-1-\frac{1}{2^{n+1}}\sum_{k=1}^{n}\frac{2^{k}}{k}\right)\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
@@
$\equiv$ $(-1)^{n}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\frac{2p+(\frac{2}{p})-P_{p}}{2p}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
@@
(5.1) $A_{q}:=\prod_{r=0}^{q-2}J_{q}(\psi,\chi_{q}^{r})=\frac{G_{q}(\psi)^{q-1}}{q}.$
(5.2) $S_{q};=\sum_{r=0}^{q-2}\frac{1}{J_{q}(\psi,\chi_{q}^{r})}=\frac{1-q}{q}\left(1+\psi(-1)\right).$
(5.3) $T_{q}:=\sum_{r=0}^{q-2}\frac{(-1)^{r}}{J_{q}(\psi,\chi_{q}^{r})}=\frac{1-q}{q}(2-\overline{\psi(2)}).$
@@
Since $\chi_{q}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$ and $\psi\neq\varepsilon$, there is an integer $j\in[1,q-2]$ such that $\chi_{q}^{-j}=\psi$. One can verify that
@@
$=(-1)^{j+1}\cdot\frac{G_{q}(\varepsilon)}{G_{q}(\chi^{-j})G_{q}(\chi^{j})}\cdot\prod_{r\in[0,q-2]}\frac{G_{q}(\chi_{q}^{-j})G_{q}(\chi_{q}^{r})}{G_{q}(\chi_{q}^{-j+r})}$
@@
$\sum_{r\in[0,q-2]}\chi_{q}^{2r}(x)=\begin{cases}(q-1)/2&\mbox{if}\ x=\pm 1,\\ 0&\mbox{otherwise.}\end{cases}$
@@
In 1987, Greene [7, Definition 2.4] used Jacobi sums to obtain the finite field analogues of binomial coefficients. In fact, for any $A,B\in\widehat{\mathbb{F}_{q}^{\times}}$, Green defined
@@
This identity can be viewed as a finite field analogue of Sury’s identity
@@
Acknowledgments The authors would like to thank Prof. Hao Pan for his helpful comments and steadfast encouragement.

### v4

@@ Abstract.
where $Q_{p}$ is the $p$-th term of the companion Pell sequence $\{Q_{i}\}_{i=0}^{\infty}$ defined by $Q_{0}=Q_{1}=2$ and $Q_{i+1}=2Q_{i}+Q_{i-1}$.
@@ 1.2. Background and motivation
Although Chapman only changed the sizes of the matrices, the calculations of $\det C_{p}^{(0)}$ and $\det C_{p}^{(1)}$ became extremely complicated. In fact, for $p\equiv 1\ ({\rm{mod}}\ 4)$, let $\epsilon_{p}>1$ and $h_{p}$ denote the fundamental unit and class number of $\mathbb{Q}(\sqrt{p})$ respectively, and write
@@
Vsemirnov [16, 17] later confirmed this tough conjecture by using ingenious and sophisticated matrix decompositions.
@@
which involves the nonzero squares over $\mathbb{F}_{p}$. Sun [13, Theorem 1.2] proved that $\det S_{p}\ {\rm mod}\ p\mathbb{Z}$ is a nonzero square over $\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}$. Moreover, Sun conjectured that $\det S_{p}$ is indeed a square of some integer whenever $p\equiv 3\ ({\rm{mod}}\ 4)$. This conjecture was later confirmed by Alekseyev and Krachun. For the case $p\equiv 1\ ({\rm{mod}}\ 4)$, if we write $p=a^{2}+4b^{2}$ with $a,b\in\mathbb{Z}$ and $a\equiv 1\ ({\rm{mod}}\ 4)$, then Cohen, Sun and Vsemirnov conjectured that $S_{p}/a$ is also a square of certain integer. This conjecture was later confirmed by the first author [18].
@@
and obtained several interesting results concerning $\det S_{p}(m)$. For example, when $p\equiv 3\ ({\rm{mod}}\ 4)$, Sun [14, Theorem 1.2(ii)] proved that
@@
$\left\{s_{1}=1,s_{2},\cdots,s_{n}\right\}=\{x^{2}:x\in\mathbb{F}_{q}^{\times}\}$
be the set of all nonzero squares over $\mathbb{F}_{q}$. Then we define the matrix
(1.5) $B_{q}(m):=\left[\left(s_{i}+s_{j}\right)^{m}\right]_{2\leq i,j\leq n}.$
@@
(1.6) $\det\left[h(x_{i}+y_{j})\right]_{1\leq i,j\leq m}=a_{m-1}^{m}\cdot\prod_{r=0}^{m-1}\binom{m-1}{r}\cdot\prod_{1\leq i<j\leq m}\left(x_{i}-x_{j}\right)\left(y_{j}-y_{i}\right).$
By this we immediately see that $B_{q}(m)$ is singular whenever $m\leq n-3$. Hence it is meaningful to consider the cases of $n-2\leq m\leq q-1$. In this paper, we focus on the cases of m=n2,n1,nm=n-2,n-1,n, and we will see later that the methods used in each case are very different.
@@ 1.3. The Pell sequence
$Q_{i}=(1+\sqrt{2})^{i}+(1-\sqrt{2})^{i}.$
@@
1.4. Main results
Now we state our first result, which involves the cases m=n1,n2m=n-1,n-2.
@@ 1.4. Main results
where h(p)h(-p) is the class number of the imaginary quadratic field $\mathbb{Q}(\sqrt{-p})$.
(ii) Suppose $p\equiv 1\ ({\rm{mod}}\ 4)$. Then
@@
Remark $1.$
With the help of a computer and by the referee’s comments, we verify that
(1.7) $\left\{p:\ p\ \text{is a prime with}\ 7\leq p\leq 10^{6}\ \text{and}\ Q_{p}\equiv 2\ ({\rm{mod}}\ p^{2}\mathbb{Z})\right\}=\left\{13,31\right\}.$
@@
We next turn to the case $m=n$.
@@
$b_{p}=\frac{2(\frac{2}{p})-2P_{p}-p}{p}\ {\rm mod}\ p\mathbb{Z}\in\mathbb{Z}/p\mathbb{Z}=\mathbb{F}_{p}.$
@@
$2P_{p}\equiv 2\left(\frac{2}{p}\right)-p\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
Remark 1.2.
As (1.7), it is natural to consider the odd prime $p$, which satisfies the congruence
$2P_{p}\equiv 2\left(\frac{2}{p}\right)-p\ ({\rm{mod}}\ p^{2}\mathbb{Z}).$
By the referee’s comments, we have
(1.8) $\left\{p:\ p\ \text{is a prime with}\ 7\leq p\leq 10^{6}\ \text{and}\ 2P_{p}\equiv 2\left(\frac{2}{p}\right)-p\ ({\rm{mod}}\ p^{2}\mathbb{Z})\right\}=\left\{29\right\}.$
@@
[…]
@@
Let $q=p^{f}\geq 3$ be an odd prime power. Then $D_{q}^{-}(\phi)$ is singular if and only if $q\equiv 3\ ({\rm{mod}}\ 4)$. For $q\equiv 1\ ({\rm{mod}}\ 4)$, we have
@@
1.5. Outline of this paper
We will prove Theorem 1.1 and its corollary in Section 2. In Section 3, we shall introduce some necessary results on almost circulant matrices and some $p$-adic tools. The proofs of Theorems 1.2 and 1.3 will be given in Section 4 and Section 5 respectively.
@@ 1.5. Outline of this paper
Recall that n=(q1)/2n=(q-1)/2 and $s_{1}=1,s_{2},\cdots,s_{n}$ are exactly all the nonzero squares over $\mathbb{F}_{q}$. Also, for any integers $a,b$ with $a\leq b$, we use the symbol $[a,b]$ to denote the set $\{a,a+1,\cdots,b\}$.
@@
We now need the following results related to some congruences involving the Pell sequence, which were obtained by Z.-H. Sun [10, Theorem 4.1], Z.-W. Sun [11, Final Remark] and Z.-W. Sun [12, Remark 3.1] respectively.
@@
where $B_{p-3}$ is the (p3)(p-3)-th Bernoulli number.
@@
$F(t)=t^{n-1}+t^{n-2}+\cdots+t+1=\prod_{2\leq j\leq n}\left(t-s_{j}\right)=\sum_{k=0}^{n-1}(-1)^{k}\sigma_{k}\left(s_{2},\cdots,s_{n}\right)t^{n-1-k},$
@@
Suppose that $f\geq 2$. As $q=p^{f}\geq 9$, we have
@@
$\det B_{q}(n-2)$ $=\prod_{r=0}^{n-2}\binom{n-2}{r}\cdot\prod_{2\leq i<j\leq n}\left(-\left(s_{j}-s_{i}\right)^{2}\right)$
@@
Proof of Corollary 1.1. (i) If $p\equiv 1\ ({\rm{mod}}\ 4)$, then $n$ is even. Thus,
@@
where h(p)h(-p) is the class number of the imaginary quadratic field $\mathbb{Q}(\sqrt{-p})$. This completes the proof of (i).
(ii) Suppose $p\equiv 1\ ({\rm{mod}}\ 4)$. Then by the Wilson congruence, one can very that
$1\equiv(p-1)!\equiv(-1)^{n}\cdot(n!)^{2}\equiv(n!)^{2}\ ({\rm{mod}}\ p\mathbb{Z}).$
@@
where the last equality follows from n(n1)=3/4n(n-1)=3/4 over $\mathbb{F}_{p}$.
Suppose now $p\equiv 3\ ({\rm{mod}}\ 4)$. Then $n$ is odd. By Theorem 1.1(ii) we clearly have
@@ 3.2. Some $p$-adic preparations
Recall that $q=p^{f}$. Let $\zeta_{q-1}\in\mathbb{C}_{p}$ be a primitive (q1)(q-1)-th root of unity. Then it is known that $\mathbb{Q}_{p}(\zeta_{q-1})/\mathbb{Q}_{p}$ is an unramified extension and $\mathbb{Q}_{p}(\zeta_{q-1}):\mathbb{Q}_{p}]=f$. Hence
@@
As $0\leq r\leq n-1$, we have $x_{0}(r)<1/2$. Hence by Eq. (3.3) we obtain
@@
Recall that $p$ is an odd prime. We next introduce the $p$-adic Gamma function $\Gamma_{p}:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}^{\times}$, where $\mathbb{Z}_{p}^{\times}$ denotes the group of all $p$-adic units. For any $n\in\mathbb{Z}^{+}$, we define
$\Gamma_{p}(n)=(-1)^{n}\prod_{k\in[1,n-1]\cap\mathbb{Z}_{p}^{\times}}k.$
@@
Let notations be as in Section 3. In this section, we fix a generator $g$ of the cyclic group $\mathbb{F}_{q}^{\times}$. Recall that n=(q1)/2n=(q-1)/2. For any $i\in[0,n-1]$, we let
@@
(4.2) $\det B_{q}(n)=\det\left[\omega_{q}^{-n}\left(1+g^{2j-2i}\right)\right]_{1\leq i,j\leq n-1}\ {\rm mod}\ \mathfrak{p}=\det W_{n}({\bm{v}})\ {\rm mod}\ \mathfrak{p},$
@@
Note first that, in this case, $\mathbb{Q}_{p}(\zeta_{q-1})=\mathbb{Q}_{p}$ and $\mathfrak{p}=p\mathbb{Z}_{p}$. By Eq. (4.5), for any $r\in[1,n-1]$, we see that
@@
We next consider $\lambda_{r}\ {\rm mod}\ p\mathbb{Z}_{p}$. Suppose $r\in[1,n-1]$. Then By Lemma 3.5 and Lemma 3.4 one can verify that
$\lambda_{r}$ $\equiv\frac{(-1)^{r}}{2}J_{p}\left(\omega_{p}^{-n},\omega_{p}^{-r}\right)$
@@
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\prod_{r=0}^{n}\binom{n}{r}\cdot\left(3-2\sum_{r=0}^{n}\binom{n}{r}^{-1}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(3-\frac{n+1}{2^{n}}\sum_{k=1}^{n+1}\frac{2^{k}}{k}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(3-\frac{1}{2^{n+1}}\sum_{k=1}^{n+1}\frac{2^{k}}{k}\right)$
$\equiv$ $(-1)^{n-1}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\left(1-\frac{1}{2^{n+1}}\sum_{k=1}^{n}\frac{2^{k}}{k}\right)\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
@@
$\equiv$ $(-1)^{n}\left(\frac{1}{2}\right)^{n-2}\cdot\frac{(n!)^{n+1}}{(0!1!\cdots n!)^{2}}\cdot\frac{2(\frac{2}{p})-2P_{p}-p}{p}\ ({\rm{mod}}\ p\mathbb{Z}_{p}).$
@@
(5.1) $A_{q}=\prod_{r=0}^{q-2}J_{q}(\psi,\chi_{q}^{r})=\frac{G_{q}(\psi)^{q-1}}{q}.$
(5.2) $S_{q}=\sum_{r=0}^{q-2}\frac{1}{J_{q}(\psi,\chi_{q}^{r})}=\frac{1-q}{q}\left(1+\psi(-1)\right).$
(5.3) $T_{q}=\sum_{r=0}^{q-2}\frac{(-1)^{r}}{J_{q}(\psi,\chi_{q}^{r})}=\frac{1-q}{q}(2-\overline{\psi(2)}).$
@@
Since $\chi_{q}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$ and $\psi\neq\varepsilon$, there is a unique integer $j\in[1,q-2]$ such that $\chi_{q}^{-j}=\psi$. One can verify that
@@
$=(-1)^{j+1}\cdot\frac{G_{q}(\varepsilon)}{G_{q}(\chi_{q}^{-j})G_{q}(\chi_{q}^{j})}\cdot\prod_{r\in[0,q-2]}\frac{G_{q}(\chi_{q}^{-j})G_{q}(\chi_{q}^{r})}{G_{q}(\chi_{q}^{-j+r})}$
@@
$\sum_{r\in[0,(q-3)/2]}\chi_{q}^{2r}(x)=\begin{cases}(q-1)/2&\mbox{if}\ x=\pm 1,\\ 0&\mbox{otherwise.}\end{cases}$
@@
In 1987, Greene [7, Definition 2.4] used Jacobi sums to obtain a finite field analogue of binomial coefficients. In fact, for any $A,B\in\widehat{\mathbb{F}_{q}^{\times}}$, Green defined
@@
The identity (5.5) can be viewed as a finite field analogue of Sury’s identity
@@
Acknowledgments The authors would like to thank the referee for helpful comments. We also thank Prof. Hao Pan for his helpful comments and steadfast encouragement.
This research was supported by the Natural Science Foundation of China (Grant Nos. 12101321, 12201291 and 12371004). The first author was supported by the Natural Science Foundation of the Higher Education Institutions of Jiangsu Province (Grant No. 25KJB110010).